OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope (-10,-1);m=-1 A.)y+10=-(x+1) B.)y+1=-(x+10) C.)y-1=-(x+10) D.)y-1=-(x-10)

OpenStudy (aykayyy):

use the formula.... y - y1 = mx - x1 ..... and your x = -10 and y = -1 I would say the answer was y+1 = -(x+10) ... B is the closest answer

OpenStudy (anonymous):

ooo ok..i c now...thank u soo much